OpenStudy (anonymous):

An ad on TV claims that 70% of doctors interviewed chose Brady's Aspirin as the leading medicine to cure minor aches. If 20 doctors were interviewed, how many chose Brady's? 10 14 15 20

OpenStudy (anonymous):

@zzr0ck3r

OpenStudy (anonymous):

@UsukiDoll

OpenStudy (usukidoll):

70/100 x 20 =?

OpenStudy (usukidoll):

70/100 <- do that first and then multiply the result by 20

OpenStudy (anonymous):

14

OpenStudy (usukidoll):

yup

OpenStudy (anonymous):

Your mom has sent you to buy a new suit, and will let you get new shoes too if you keep it under $200. The suit you want to buy is $273. If you are getting 20% off the price, will you get the shoes? Yes, the suit will cost less than $200. No, the suit will cost more than $200. The cost will be exactly $200. Not enough information.

OpenStudy (anonymous):

what about this one

OpenStudy (usukidoll):

20/100 x 273 = 54.60 273-54.60 = 218.40 can't get the shoes.

OpenStudy (anonymous):

so its not enough information

OpenStudy (usukidoll):

we do have enough information. after applying the discount, I'm still overbudget

OpenStudy (anonymous):

so it b No, the suit will cost more than $200.

OpenStudy (usukidoll):

yes the suit is still over $200

OpenStudy (usukidoll):

second choice is what you need no the suit is more than $200

OpenStudy (anonymous):

i only need help on a couple moe then im done

OpenStudy (usukidoll):

I'll try.